# Benchmark stars for mean stellar density and surface gravity estimates of solar-type stars

### v1

@@ Appendix A Data used for re-calibration of empirical mass and radius polynomial functions
47 Tuc E32 40.913 $-0.71$ $-0.0646$ $0.0731$ 3.7800 $-0.0826$ $0.0019$ 3.7750
47 Tuc V69 29.540 $-0.71$ $-0.0574$ $0.1189$ 3.7750 $-0.0661$ $0.0651$ 3.7770
Kepler-35 20.734 $-0.34$ $-0.0517$ $0.0122$ 3.7490 $-0.0918$ $-0.1045$ 3.7160
RW Lac 10.369 $-0.30$ $-0.0325$ $0.0741$ 3.7600 $-0.0605$ $-0.0159$ 3.7450
KIC 6131659 17.528 $-0.23$ $-0.0289$ $-0.0533$ 3.7528 $-0.1551$ $-0.1859$ 3.6794
Kepler-453 27.322 $+0.09$ $-0.0250$ $-0.0794$ 3.7420
TYC 5962-2159-1 8.220 $+0.09$ $-0.0195$ $-0.0013$ 3.7400 $-0.1713$ $-0.1612$ 3.5990
Kepler-47 7.448 $-0.25$ $-0.0191$ $-0.0287$ 3.7510
V565 Lyr 18.799 $+0.28$ $-0.0020$ $0.0418$ 3.7480 $-0.0318$ $-0.0129$ 3.7350
V530 Ori 6.111 $-0.12$ $0.0016$ $-0.0088$ 3.7770
V818 Tau 5.609 $+0.05$ $0.0105$ $-0.0348$ 3.7520 $-0.1292$ $-0.1314$ 3.6330
KIC 7177553 17.996 $-0.05$ $0.0183$ $-0.0269$ 3.7630 $-0.0061$ $-0.0264$ 3.7590
Kepler-34 27.796 $-0.07$ $0.0203$ $0.0651$ 3.7720 $0.0089$ $0.0385$ 3.7680
TIC 71877648 32.394 $+0.09$ $0.0199$ $0.3462$ 3.7150
TYC 6296-96-1 6.527 $+0.07$ $0.0327$ $0.0233$ 3.7730 $0.0278$ $0.0179$ 3.7690
V963 Cen 15.269 $-0.06$ $0.0339$ $0.1600$ 3.7640 $0.0315$ $0.1527$ 3.7650
V568 Lyr 14.470 $+0.28$ $0.0361$ $0.1452$ 3.7520 $-0.0822$ $-0.1072$ 3.6830
TYC 1243-402-1 8.074 $+0.00$ $0.0374$ $0.0969$ 3.7750 $-0.0177$ $-0.0269$ 3.7450
QR Hya 5.006 $-0.01$ $0.0415$ $0.1087$ 3.7730 $0.0283$ $0.0802$ 3.7670
AL Dor 14.905 $-0.10$ $0.0421$ $0.0381$ 3.7820 $0.0425$ $0.0407$ 3.7820
V785 Cep 6.504 $-0.06$ $0.0426$ $0.1535$ 3.7710 $0.0338$ $0.1377$ 3.7690
EPIC 211409263 69.728 $+0.03$ $0.0457$ $0.0298$ 3.7850 $-0.1261$ $-0.1469$ 3.6730
TYC 6296-2012-1 11.991 $+0.07$ $0.0496$ $0.0711$ 3.7880 $-0.1349$ $-0.1938$ 3.6530
KX Cnc 31.220 $+0.07$ $0.0548$ $0.0223$ 3.7710 $0.0508$ $0.0250$ 3.7670
HD 219869 6.062 $-0.39$ $0.0554$ $0.2148$ 3.7970 $0.0026$ $0.0453$ 3.7950
TYC 7091-888-1 11.658 $-0.33$ $0.0624$ $0.2634$ 3.8010 $-0.1061$ $-0.1372$ 3.7320
KIC 7037405 207.108 $-0.27$ $0.0453$ $0.2420$ 3.7850
EW Ori 6.937 $+0.05$ $0.0693$ $0.0674$ 3.7830 $0.0504$ $0.0402$ 3.7710
KIC 9970396 235.299 $-0.35$ $0.0013$ $0.0449$ 3.7940
LV Her 18.436 $+0.08$ $0.0766$ $0.1329$ 3.7930 $0.0681$ $0.1183$ 3.7800
AI Phe 24.592 $-0.14$ $0.0769$ $0.2565$ 3.7920 $0.0948$ $0.4673$ 3.7070
LL Aqr 20.178 $+0.02$ $0.0777$ $0.1209$ 3.7840 $0.0144$ $0.0009$ 3.7560
Kepler-1647 11.259 $-0.14$ $0.0866$ $0.2529$ 3.7930 $-0.0142$ $-0.0149$ 3.7610
KIC 3439031 5.952 $+0.10$ $0.0884$ $0.1484$ 3.8150 $0.0890$ $0.1471$ 3.8150
TIC 172900988 19.658 $+0.34$ $0.0926$ $0.1407$ 3.7820 $0.0799$ $0.1186$ 3.7770
KIC 8430105 63.327 $-0.46$ $-0.0897$ $-0.1253$ 3.7560
CQ Ind 8.974 $-0.04$ $0.1036$ $0.1491$ 3.8100 $0.0528$ $0.0518$ 3.7880
KIC 7821010 24.238 $+0.10$ $0.1062$ $0.1059$ 3.8270 $0.0867$ $0.0828$ 3.8170
V338 Vir 5.985 $-0.10$ $0.1164$ $0.1855$ 3.8180
CPD -54 810 26.132 $+0.00$ $0.1167$ $0.2853$ 3.8100 $0.0373$ $0.0724$ 3.8010
FM Leo 6.729 $-0.11$ $0.1187$ $0.2107$ 3.8060 $0.1039$ $0.1785$ 3.8050
ASAS J085002+1752.5 5.226 $-0.01$ $0.1225$ $0.1962$ 3.8170
LX Mus 11.751 $+0.09$ $0.1282$ $0.1287$ 3.8150 $0.1317$ $0.1449$ 3.8170
HD 22064 9.135 $-0.05$ $0.1287$ $0.1915$ 3.8300
RU Cnc 10.173 $-0.26$ $0.1294$ $0.3400$ 3.8170
BW Aqr 6.720 $-0.07$ $0.1377$ $0.2388$ 3.8100
KIC 9540226 175.443 $-0.31$ $0.0065$ $0.0145$ 3.7650
BK Peg 5.490 $-0.12$ $0.0993$ $0.1685$ 3.8010
HD 149946 23.310 $-0.19$ $0.1526$ $0.3339$ 3.8220 $0.0573$ $0.0878$ 3.8190
ASAS J090232-5653.4 20.822 $-0.15$ $0.1572$ $0.1838$ 3.8470 $0.1374$ $0.1436$ 3.8250
KIC 8410637 408.324 $+0.02$ $0.1169$ $0.1920$ 3.8048
KIC 5640750 987.398 $-0.29$ $0.1113$ $0.2679$ 3.7818
HD 32129 16.412 $+0.19$ $0.1872$ $0.2724$ 3.8270 $0.0289$ $-0.0055$ 3.7600
KIC 11285625 10.790 $-0.58$ $0.1884$ $0.3269$ 3.8430 $0.0792$ $0.1679$ 3.8570
KIC 9777062 19.230 $-0.03$ $0.1520$ $0.1886$ 3.8540
V501 Mon 7.021 $+0.01$ $0.1640$ $0.2019$ 3.8450
BD +37 410 15.535 $-0.03$ $0.2348$ $0.4622$ 3.8210 $0.0700$ $0.0741$ 3.8010
HR 2214 23.810 $-0.23$ $0.2475$ $0.3054$ 3.8510
TZ For 75.666 $+0.02$ $0.2918$ $0.5988$ 3.8030
@@ Appendix B Tables of results for transiting exoplanet systems
[…]
@@
[…]
@@ Appendix C KIC 6131659
$M_{1}$ [$M_{\odot}$] $0.9355\pm 0.0036$
$M_{2}$ [$M_{\odot}$] $0.6997\pm 0.0023$
$R_{1}$ [$R_{\odot}$] $0.8844\pm 0.0012$
$R_{2}$ [$R_{\odot}$] $0.6518\pm 0.0020$

### v2

@@ Appendix A Data used for re-calibration of empirical mass and radius polynomial functions
47 Tuc E32 40.913 0.71-0.71 0.0646-0.0646 $0.0731$ 3.7800 0.0826-0.0826 $0.0019$ 3.7750
47 Tuc V69 29.540 0.71-0.71 0.0574-0.0574 $0.1189$ 3.7750 0.0661-0.0661 $0.0651$ 3.7770
Kepler-35 20.734 0.34-0.34 0.0517-0.0517 $0.0122$ 3.7490 0.0918-0.0918 0.1045-0.1045 3.7160
RW Lac 10.369 0.30-0.30 0.0325-0.0325 $0.0741$ 3.7600 0.0605-0.0605 0.0159-0.0159 3.7450
KIC 6131659 17.528 0.23-0.23 0.0289-0.0289 0.0533-0.0533 3.7528 0.1551-0.1551 0.1859-0.1859 3.6794
Kepler-453 27.322 +0.09+0.09 0.0250-0.0250 0.0794-0.0794 3.7420
TYC 5962-2159-1 8.220 +0.09+0.09 0.0195-0.0195 0.0013-0.0013 3.7400 0.1713-0.1713 0.1612-0.1612 3.5990
Kepler-47 7.448 0.25-0.25 0.0191-0.0191 0.0287-0.0287 3.7510
V565 Lyr 18.799 +0.28+0.28 0.0020-0.0020 $0.0418$ 3.7480 0.0318-0.0318 0.0129-0.0129 3.7350
V530 Ori 6.111 0.12-0.12 $0.0016$ 0.0088-0.0088 3.7770
V818 Tau 5.609 +0.05+0.05 $0.0105$ 0.0348-0.0348 3.7520 0.1292-0.1292 0.1314-0.1314 3.6330
KIC 7177553 17.996 0.05-0.05 $0.0183$ 0.0269-0.0269 3.7630 0.0061-0.0061 0.0264-0.0264 3.7590
Kepler-34 27.796 0.07-0.07 $0.0203$ $0.0651$ 3.7720 $0.0089$ $0.0385$ 3.7680
TIC 71877648 32.394 +0.09+0.09 $0.0199$ $0.3462$ 3.7150
TYC 6296-96-1 6.527 +0.07+0.07 $0.0327$ $0.0233$ 3.7730 $0.0278$ $0.0179$ 3.7690
V963 Cen 15.269 0.06-0.06 $0.0339$ $0.1600$ 3.7640 $0.0315$ $0.1527$ 3.7650
V568 Lyr 14.470 +0.28+0.28 $0.0361$ $0.1452$ 3.7520 0.0822-0.0822 0.1072-0.1072 3.6830
TYC 1243-402-1 8.074 +0.00+0.00 $0.0374$ $0.0969$ 3.7750 0.0177-0.0177 0.0269-0.0269 3.7450
QR Hya 5.006 0.01-0.01 $0.0415$ $0.1087$ 3.7730 $0.0283$ $0.0802$ 3.7670
AL Dor 14.905 0.10-0.10 $0.0421$ $0.0381$ 3.7820 $0.0425$ $0.0407$ 3.7820
V785 Cep 6.504 0.06-0.06 $0.0426$ $0.1535$ 3.7710 $0.0338$ $0.1377$ 3.7690
EPIC 211409263 69.728 +0.03+0.03 $0.0457$ $0.0298$ 3.7850 0.1261-0.1261 0.1469-0.1469 3.6730
TYC 6296-2012-1 11.991 +0.07+0.07 $0.0496$ $0.0711$ 3.7880 0.1349-0.1349 0.1938-0.1938 3.6530
KX Cnc 31.220 +0.07+0.07 $0.0548$ $0.0223$ 3.7710 $0.0508$ $0.0250$ 3.7670
HD 219869 6.062 0.39-0.39 $0.0554$ $0.2148$ 3.7970 $0.0026$ $0.0453$ 3.7950
TYC 7091-888-1 11.658 0.33-0.33 $0.0624$ $0.2634$ 3.8010 0.1061-0.1061 0.1372-0.1372 3.7320
KIC 7037405 207.108 0.27-0.27 $0.0453$ $0.2420$ 3.7850
EW Ori 6.937 +0.05+0.05 $0.0693$ $0.0674$ 3.7830 $0.0504$ $0.0402$ 3.7710
KIC 9970396 235.299 0.35-0.35 $0.0013$ $0.0449$ 3.7940
LV Her 18.436 +0.08+0.08 $0.0766$ $0.1329$ 3.7930 $0.0681$ $0.1183$ 3.7800
AI Phe 24.592 0.14-0.14 $0.0769$ $0.2565$ 3.7920 $0.0948$ $0.4673$ 3.7070
LL Aqr 20.178 +0.02+0.02 $0.0777$ $0.1209$ 3.7840 $0.0144$ $0.0009$ 3.7560
Kepler-1647 11.259 0.14-0.14 $0.0866$ $0.2529$ 3.7930 0.0142-0.0142 0.0149-0.0149 3.7610
KIC 3439031 5.952 +0.10+0.10 $0.0884$ $0.1484$ 3.8150 $0.0890$ $0.1471$ 3.8150
TIC 172900988 19.658 +0.34+0.34 $0.0926$ $0.1407$ 3.7820 $0.0799$ $0.1186$ 3.7770
KIC 8430105 63.327 0.46-0.46 0.0897-0.0897 0.1253-0.1253 3.7560
CQ Ind 8.974 0.04-0.04 $0.1036$ $0.1491$ 3.8100 $0.0528$ $0.0518$ 3.7880
KIC 7821010 24.238 +0.10+0.10 $0.1062$ $0.1059$ 3.8270 $0.0867$ $0.0828$ 3.8170
V338 Vir 5.985 0.10-0.10 $0.1164$ $0.1855$ 3.8180
CPD -54 810 26.132 +0.00+0.00 $0.1167$ $0.2853$ 3.8100 $0.0373$ $0.0724$ 3.8010
FM Leo 6.729 0.11-0.11 $0.1187$ $0.2107$ 3.8060 $0.1039$ $0.1785$ 3.8050
ASAS J085002+1752.5 5.226 0.01-0.01 $0.1225$ $0.1962$ 3.8170
LX Mus 11.751 +0.09+0.09 $0.1282$ $0.1287$ 3.8150 $0.1317$ $0.1449$ 3.8170
HD 22064 9.135 0.05-0.05 $0.1287$ $0.1915$ 3.8300
RU Cnc 10.173 0.26-0.26 $0.1294$ $0.3400$ 3.8170
BW Aqr 6.720 0.07-0.07 $0.1377$ $0.2388$ 3.8100
KIC 9540226 175.443 0.31-0.31 $0.0065$ $0.0145$ 3.7650
BK Peg 5.490 0.12-0.12 $0.0993$ $0.1685$ 3.8010
HD 149946 23.310 0.19-0.19 $0.1526$ $0.3339$ 3.8220 $0.0573$ $0.0878$ 3.8190
ASAS J090232-5653.4 20.822 0.15-0.15 $0.1572$ $0.1838$ 3.8470 $0.1374$ $0.1436$ 3.8250
KIC 8410637 408.324 +0.02+0.02 $0.1169$ $0.1920$ 3.8048
KIC 5640750 987.398 0.29-0.29 $0.1113$ $0.2679$ 3.7818
HD 32129 16.412 +0.19+0.19 $0.1872$ $0.2724$ 3.8270 $0.0289$ 0.0055-0.0055 3.7600
KIC 11285625 10.790 0.58-0.58 $0.1884$ $0.3269$ 3.8430 $0.0792$ $0.1679$ 3.8570
KIC 9777062 19.230 0.03-0.03 $0.1520$ $0.1886$ 3.8540
V501 Mon 7.021 +0.01+0.01 $0.1640$ $0.2019$ 3.8450
BD +37 410 15.535 0.03-0.03 $0.2348$ $0.4622$ 3.8210 $0.0700$ $0.0741$ 3.8010
HR 2214 23.810 0.23-0.23 $0.2475$ $0.3054$ 3.8510
TZ For 75.666 +0.02+0.02 $0.2918$ $0.5988$ 3.8030
@@ Appendix B Tables of results for transiting exoplanet systems
Table 6: Results from the analysis of the Keper light curve of HAT-P-7. $T_{0}$ is the Julian date of the barycentric dynamical time of mid transit (${\rm BJD}_{\rm TDB}$). The number of observations used in the analysis for each data set are labelled $N_{\it Kepler}$, $N_{\rm rv,1}$, etc. The root mean square residual (rms) of the residuals from the maximum-likelihood model fit for each data set are labelled $\sigma_{\it Kepler}$, $\sigma_{\rm rv,1}$, etc.
@@
Table 9: Results from the analysis of the Kepler light curve of Kepler-1. $T_{0}$ is the Julian date of the barycentric dynamical time of mid transit (${\rm BJD}_{\rm TDB}$). The number of observations used in the analysis for each data set are labelled $N_{\it Kepler}$, $N_{\rm rv,1}$, etc. The root mean square residual (rms) of the residuals from the maximum-likelihood model fit for each data set are labelled $\sigma_{\it Kepler}$, $\sigma_{\rm rv,1}$, etc.
@@ Appendix C KIC 6131659
$_{1} [MM_{\odot}$] $0.9355\pm 0.0036$
$_{2} [MM_{\odot}$] $0.6997\pm 0.0023$
$_{1} [RR_{\odot}$] $0.8844\pm 0.0012$
$_{2} [RR_{\odot}$] $0.6518\pm 0.0020$